OpenStudy (anonymous):

I WILL GIVE MEDAL AND BECOME A FAN!!!!!! Salisha purchased 45¢ and 14¢ stamps at the post office. She bought 34 stamps in all for $10.96. How many of each type of stamp did she purchase? Show both equations (one for the total of the stamps and one for the value of the stamps) Then solve using substitution.

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

[ffff00] my name is jeff lel ok lets see

OpenStudy (anonymous):

@catlikesu

OpenStudy (anonymous):

oh what no color codes

OpenStudy (anonymous):

@crazzydave

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

umm tough one man

OpenStudy (anonymous):

ikr imao

OpenStudy (sleepyjess):

We need to create a system of equations

OpenStudy (anonymous):

i aggree

OpenStudy (mathstudent55):

Let x = number of $0.45 stamps. Let y = number of $0.14 stamps. Using x and y, how do you show the total number of stamps bought?

OpenStudy (sleepyjess):

45 cent stamps are x 14 cent stamps are y We know that she bought 34 stamps

OpenStudy (anonymous):

lel let @sleepyjess handel this imao

OpenStudy (anonymous):

ok so what do i have now

OpenStudy (anonymous):

?

OpenStudy (sleepyjess):

Sorry, was afk

OpenStudy (anonymous):

can you help

OpenStudy (sleepyjess):

What do you think the first equation will be?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

If you aren't willing to try, then I'm not willing to help

OpenStudy (anonymous):

if Salisha Brought 34 Stamps Total And It Came Out To 10.96 It Means She Would Have Brought 10.96 dollars total. the equation to show you the amount of the stamps is -----> Let x = number of $0.45 stamps. Let y = number of $0.14 stamps. The Next Equation to the total of stamps

OpenStudy (anonymous):

is that good

OpenStudy (sleepyjess):

Well, what is the equation

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

y*x= 34

OpenStudy (sleepyjess):

good try, but it would actually be x+y = 34

OpenStudy (sleepyjess):

Now, we know that x = .45 and y = .14

OpenStudy (anonymous):

ok the second one i said is x*y=10.96

OpenStudy (sleepyjess):

Nope, no multiplication yet

OpenStudy (anonymous):

ok so subtraction

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

\(\color{blue}{\text{Originally Posted by}}\) @jazzy288a I WILL GIVE MEDAL AND BECOME A FAN!!!!!! Salisha purchased 45¢ and 14¢ stamps at the post office. She bought 34 stamps in all for $10.96. How many of each type of stamp did she purchase? Show both equations (one for the total of the stamps and one for the value of the stamps) Then solve using substitution. \(\color{blue}{\text{End of Quote}}\) You already have one equation... and that is: \(\sf x+y=34\) where x is the 45 cents and y is the 15 cent stamp.

TheSmartOne (thesmartone):

Can you make another equation with this given information? "Salisha purchased 45¢ and 14¢ stamps at the post office." and "all for $10.96"

OpenStudy (anonymous):

x-y=10.96

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

that is wrong...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do i divide

TheSmartOne (thesmartone):

Ok, anyways. I don't got all day. Your second equation is: \(\sf.45x+.14y=10.96\)

TheSmartOne (thesmartone):

\(\sf x+y=34\) \(\sf.45x+.14y=10.96\) Solve the system of equations...

OpenStudy (anonymous):

is that it thanks so much

TheSmartOne (thesmartone):

Solve that system of equations using substituition. So what I just gave you isn't the answer...

OpenStudy (mathstudent55):

@jazzy288a Here is the explanation for the two equations. Let x = the number of 45 cent stamps Let y = number of 14 cent stamps 1. The first equation deals with the number of stamps that were bought The total number of stamps bought is the number of 45-cent stamps plus the number of 14-cent stamps. Since those numbers are represented by x and y, respectively, the total number of stamps bought is x + y We are told that altogether, 34 stamps were bought, so x + y = 34 That is our first equation.

OpenStudy (mathstudent55):

2. The second equation deals with the prices of the stamps and the total amount of money spent. There were x stamps bought at 45 cents each. There were y stamps bought at 14 cents each. In dollars, 45 cents is 0.45, and in dollars, 14 cents is 0.14. x number of $0.45 stamps cost 0.45x. y number of $0.14 stamps cost 0.14y. The total amount of money spent is 0.45x + 0.14y We are told the total amount of money spent is 10.96, so we have our second equation: 0.45x + 0.14y = 10.96 This is our second equation.

OpenStudy (mathstudent55):

The two equations are: x + y = 34 0.45x + 0.14y = 10.96

OpenStudy (mathstudent55):

For the next part, you need to solve the system of equations sing substitution. To do that, solve the first equation for x. Then substitute that into the x variable of the second equation. You will have an equation in only x that can be solved. Solve for x. That tells you how many 45-cent stamps were bought. Then subtract that number from 34 to find how many 14-cent stamps were bought.